Project

General

Profile

Actions

RunPlan20151124

Expert run. Aim is to shakedown systems for December run & check through tof calibrations.

Daily schedule

November 24
estimated time 8am - 4pm (8pm)
MOM E. Overton
Shifter S. Boyd
Shifter P. Kyberd
BLOC H. Nebrensky
SOC D. Rajaram
Aim Test Run
Beam? Yes
Additional Personnel Experts

Plan & Personell:

  • 8.30am:
    Search Hall (MOM + Shifter1)
    Start setting up beamline (BLOC).
  • 9.00 am:
    Hall Searched.
    Setup for pion reference beam.
    Start DAQ when ready
    Check that run control is working with cdb (and cdb is being written)
  • 10.00am:
    Check the online reconstruction / monitoring plots for good data.
    Check alarm handlers.
    Speak to physics shifter to ensure they are also getting data / it looks OK?
  • 11.00am:
    Switch to running TOF calibration beam.
  • 11.15am (MOM, BLOC, DC goto unrelated Meeting).
  • 12.30:
    Shifter2 arrives fed to help relieve shifter1/experts for lunch.
  • 16.00:
    If everything goes well we stop taking data..

Otherwise, we have Shifter2 to help an expert debug problems until 20.00.

I think this is a plan... its a test run, so it will probably get heavily modified on the day.

Data Requested

  • Poon Reference
  • TOF Calibration

Beamline Configurations

We expect to be running with 2V beamloss and beam bump active.

Reference Run without Decay Solenoid

Particle Species p at Tgt p@D1 p@D2 p@Tof0 Proton Absorber Q1 Q2 Q3 D1 DS D2 Q4 Q5 Q6 Q7 Q8 Q9 Ran? Similar Run(s)
MeV/c MeV/c MeV/c MeV/c mm A A A A A A A A A A A A DAQ run #
pion 272.42 268.67 265.98 143.4 29 57.8 105.6 64.6 202.8 ---- 100.4 176.8 237.1 157.2 157.6 238.5 203.7

Summary of run:

We completed the data taking with success, however there are a number of points where we can improve upon for December. These are outlined below:

Controls:
  • KL monitor GUI is blank, but voltages can be read using normal SY527 control GUI.
  • Alarm Handler documentation (has been updated - thanks Paolo.).
Detectors:
  • EMR was cabled for LED running - now switched to beam mode. Does this change need documenting?
Run Control:
  • Initially was operated in expert only mode, since cooling channel was recently added. Needs verifying & documenting for Dec running.
Online Monitoring:
  • EPICS alarm work (a little too well). Quite a lot of false positives caused by empty spills (from 0 triggers), alarm thresholds need tuning.
  • EMR plots are empty.
  • Would be helpful to have more unpacking diagnostics:
    • What detector is causing the unpacking errors (from geo id).
    • In the event of 0 triggers, say that it is consistent.
Online Reconstruction:
  • Links to documentation should be provided from the webpage.
  • Would be nice to have a "help" or descriptions added.
  • Run # should be visible on plots.
    More comments in OnRec_Plots.pdf
Physics Shifter:
  • Physics devil was run
  • bugs were found which need fixing.
  • data transfer to the devil needs automating.

Updated by Overton, Edward almost 8 years ago ยท 6 revisions